# Bosonization in $R$-paraparticle Luttinger models

### v1

@@ 1 Introduction
Outside anyons, multiple theories of exotic particles that do not follow bosonic or fermionic statistics have also been proposed: all of which have not yet been observed experimentally. Probably the most well-studied are Green’s paraparticles [5], which to this day garners opposing opinions on its actual existence in nature [6, 7, 8]. While superselection rules [9] and no-go theorems [10, 11] tell us that under certain assumptions Green’s paraparticles can always be re-expressed in terms of ordinary bosons and fermions [12, 13], these studies do not explicitly rule out the possibility of them existing in nature. The field has become purely theoretical, with paraparticles being explored through their thermodynamic signatures [14], using para-manifolds [15], and as potential candidates for dark matter [16, 17, 18]. To date, the closest we have been to observing paraparticles are signatures of para-bose and para-fermi oscillators in trapped ions [19, 20, 21, 22].
Recently, an alternative formulation was proposed which makes use of the generalized commutation relations due to a four-tensor $R$ [23]. In this “$R$-paraparticle” formulation, quasiparticle excitations with parastatistical behavior can be constructed using nonlocal string operators from an underlying set of local spin operators. This approach evades the no-go theorems [10, 11] which assumes that excitations are only created by local operators. Moreover, the superselection rules [9] are inherently satisfied by quasiparticle excitations. This construction from local spin operators opens the possibility of realizing emergent paraparticles in condensed matter systems and in quantum simulations using trapped ions [24]. However, the discussions in [23] primarily focus on free paraparticle systems. In contrast, conventional bose and fermi systems exhibit a wealth of collective excitations and emergent phases arising from particle interactions. This naturally raises the question of how interactions might influence the behavior of paraparticles, making it an intriguing direction for further investigation.
@@
In this paper, we study a generalized version of the Luttinger model where we extend the conventional fermionic operators to paraparticle operators. This approach of extending the operators to paraparticle operators has recently been explored in the context of para-SYK models [36, 37]. We show that bosonization also occurs for $R$-paraparticles with fermi-like average occupation numbers [14], which we simply call “parafermions” in analogy to Green’s paraparticles [5, 9]. This presents a method in which interacting paraparticle systems can be probed using standard methods of solving bosonic systems [29]. Before proceeding to the remainder of the study, we first clarify that the term “parafermion” should not be confused with generalizations of Majorana fermions [38, 39, 40], which unfortunately share the same terminology in condensed matter literature. For instance, the work in [41] also addresses the bosonization of parafermions; however, the parafermions considered there are those related to Majorana fermions and are not the R-type parafermions considered in this work.
The paper shall be organized as follows. In Sec. 2, we review the $R$-paraparticle formulation and clarify how parabosons and parafermions are classified based on their thermodynamics and exclusion statistics. We show that $R$-parafermions obey Pauli’s exclusion principle or one of its generalizations, prompting us to define $p$-orderedness in the $R$-paraparticle formalism. In Sec. 3, parafermionic density wave excitations are shown to be bosonic, independent of the number of degrees of freedom of the paraparticle, and are completely decoupled to flavor waves. However, for systems with two internal degrees of freedom, flavor waves, which are generalizations to spin waves, are bosonic only for specific parafermion species. More importantly, we show that bosonization only applies at low energies for paraparticles that satisfy Pauli exclusion principle, as $R$-parafermions obeying higher-order exclusion principles generate partition functions that do not match in the parafermionic and bosonic bases. These results point to possible experimental realizations of paraparticle signatures in one-dimensional systems by searching for signs of flavor-charge separation [31] in the absence of magnon excitations [42], as will be discussed in Sec. 4. Finally, we summarize the contents of this paper and provide possible extensions of this work in Sec. 5.
@@ 2.1 Formulation
where $x\equiv e^{-\beta\epsilon}$, $\beta=1/k_{B}T$, and $\epsilon$ is the paraparticle energy. $d_{n}$ are nonnegative integers and describe the dimension of the $n$-particle state. Importantly, $d_{n}=0$ implies that the $n$ paraparticles cannot occupy the same mode at once. The average occupation number then can be calculated as
@@ 2.2 R-parafermions
Analogous to Green’s formulation which generalizes parafermions with integer- and half-integer spins as parabosons and parafermions [5], we also construct a classification scheme for $R$-paraparticles. Suppose the GCR due to $R$ allows at most $n^{\prime}$ particles to occupy a single mode. This implies that $d_{n>n^{\prime}}=0$, hence the average occupation number is reduced to
@@
i.e., the occupation number close to absolute zero is proportional to a Heaviside step function $\theta(x)$, and therefore the $R$-paraparticle possesses a fermi surface structure. This motivates us to classify paraparticles that satisfy Eq. (11b) as $R$-parafermions. Analogously, we show in the Appendix that when $d_{n}\neq 0$ for all $n$, the average occupation of the particle resembles bose statistics, and therefore these $R$-paraparticles are categorized as $R$-parabosons. In Green’s paraparticle formulation, these behaviors were also observed in [14].
The requirement that $d_{n>n^{\prime}}=0$ implies an exclusion statistics is at play within the $R$-paraparticle’s GCR. In general, this can be expressed as
$(\hat{\psi}_{i,a}^{+})^{p+1}=0\quad\text{while}\quad(\hat{\psi}_{i,a}^{+})^{p}\neq 0,$ (12)
where $p=1$ is the usual Pauli exclusion principle. This definition is analogous to Green’s $p$-ordered parafermions, to which there can be at most $p$ particles in the same quantum state [9]. We therefore define a $p$-ordered $R$-parafermions as paraparticles that satisfy Eq. (12).
@@
where $\alpha\in\mathbb{C}_{\neq 0}$. Ordinary fermions are retrieved by $\alpha=-1$ in Eq. (13b). Likewise, $p=2$ cases also exist for $R$-paraparticles with $m=2$, such as those described by
@@
also with $\beta\in\mathbb{C}_{\neq 0}$. The exclusion statistics of these particles go as $\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,b}^{+}=-\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,b}^{+}=0$ and $\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,a}^{+}\propto\hat{\psi}_{i,b}^{+}\hat{\psi}_{i,b}^{+}\neq 0$ for $\in\{1,2\}, aba\neq b$. This results to $\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,a}^{+}\propto\allowbreak\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,b}^{+}\hat{\psi}_{i,b}^{+}=0$, implying that three particles cannot occupy the same mode at once, and therefore $d_{n>2}=0$.
3 Parafermion bosonization
Having defined $p$-ordering and the classification of paraparticles into parafermions or parabosons, let us now consider the bosonization of one-dimensional interacting paraparticles. We define the paraparticle Luttinger model kinetic Hamiltonian term as
@@ 3 Parafermion bosonization
where the constant $\theta(rk-k_{F})$ imposes the normal ordering denoted by $\ldots:$, and $\pm k_{F}$ are the parafermi points of the 1D system. The indices $k$ and $a$ denote momentum and flavor, while $r\in\{+,-\}$ (a mode index) differentiates between the two branches of the dispersion relation $\varepsilon_{r}(k)=v_{F}(rk-k_{F})$, with $v_{F}$ as the fermi velocity. In the Luttinger model, the total density operator can be decomposed in terms of the left-moving and right-moving parafermions, $\hat{\rho}(k)=\hat{\rho}_{+}(k)+\hat{\rho}_{-}(k)$, where
@@ 3.1 Density wave bosonization and flavor-charge separation
which then imply that the density waves of all parafermionic Luttinger models have bosonic properties.
@@ 3.2 Flavor wave bosonization
$\hat{\sigma}_{r}(-q),\hat{\sigma}_{r^{\prime}}(q^{\prime})]=\sum_{k,k^{\prime},a,a^{\prime}}\alpha_{a}\alpha_{a^{\prime}}[\hat{\psi}^{+}_{r,k-q,a}\hat{\psi}_{r,k,a},\hat{\psi}^{+}_{r^{\prime},k^{\prime}+q^{\prime},a^{\prime}}\hat{\psi}_{r^{\prime},k^{\prime},a^{\prime}}]$ (20)
explicitly for specific $R^{ab}_{cd}$. In this case, bosonic flavor waves occur only if the $R$-tensor satisfies
@@
to which Eq. (20) becomes
@@
allows for the establishment of $\hat{c}_{k},\hat{c}^{+}_{k}]=\delta_{kk^{\prime}}$. $R$-tensors in the form of Eq. (13a), (13b), and (13d) with $\beta^{2}=1$ all follow these results. However, parafermions described by Eqs. (13c) and (13d) with $\beta^{2}\neq 1$ do not satisfy Eq. (21) and therefore their parafermionic Luttinger models do not have bosonic flavor wave profiles.
@@ 3.3 Equivalence spectrum
We now check wether the system described using parafermion operators are the same as that described by bosonic ones. For $m=2$ $R$-parafermions, it can be shown that the free Luttinger model Hamiltonian obeys the commutation relation
@@
where $\hat{n}_{k}=\hat{b}_{k}^{+}\hat{b}_{k}$ and $\hat{N}_{r,a}=\sum_{k}(\hat{\psi}_{r,k,a}^{+}\hat{\psi}_{r,k,a}-\theta(rk-k_{F}))$. Note that these results are applicable only to parafermions with bosonic flavor waves. The second summation term takes care of the normal ordering imposed in the parafermion operator basis. The flavor wave contribution to the Hamiltonian is the same as Eq. (25), but with $\hat{b}_{k}^{+}\hat{b}_{k}$ replaced by $\hat{c}_{k}^{+}\hat{c}_{k}$ [29], hence we consider the total Hamiltonian
@@
where $H_{0}$ is denoted by Eqs. (14) and (25) in the parafermion and boson operator bases, respectively.
One way to verify the correspondence between these Hamiltonian forms is to show that their degeneracies match in both bases. The case where parafermions have no internal degrees of freedom (i.e., spinless fermions) has already been done in [26] by evaluating the partition functions of each basis separately. Following this method, the parafermion partition function is calculated as
@@
where $y\equiv e^{-\beta v_{F}\pi/L}$. The (2n1)(2n-1) exponent is obtained by assuming that the system is under periodic boundary conditions such that $k=2n\pi/L$ (where $n$ are energy level indices) and $k_{F}=n_{F}\pi/L$. The variable $p$ denotes the order of the parafermion as discussed in Sec. 2.2. In particular, for $p=1$, we have
@@
Comparison between Eqs. (28a) and (29) shows that in the low-temperature limit (i.e., $\beta\to\infty$, and subsequently $y\to 0$), the boson and $p=1$ $R$-parafermion partition functions coincide, hence the spectrum of the free Luttinger model is equivalent in the paraparticle and boson operator basis only if the parafermion obeys Pauli’s exclusion principle. This result generalizes the idea that the Luttinger model is acceptable only at low-energies [46]. On the other hand, if the $R$-parafermion possesses a higher-order exclusion principle (such as in the case of $p=2$ in Eq. (28b)), $z_{PF}$ quickly diverges. This means that bosonization is not applicable for $R$-parafermions with $p>1$, and possibly the Luttinger model is not appropriate for modelling such systems.

### v2

@@ 1 Introduction
Outside anyons, multiple theories of exotic particles that do not follow bosonic or fermionic statistics have also been proposed: all of which have not yet been observed experimentally. Probably the most well-studied are Green’s paraparticles [5], which to this day garners opposing opinions on its actual existence in nature [6, 7, 8]. While superselection rules [9] and no-go theorems [10, 11] tell us that under certain assumptions Green’s paraparticles can always be re-expressed in terms of ordinary bosons and fermions [12, 13], these studies do not explicitly rule out the possibility of them existing in nature. The field has become purely theoretical, with paraparticles being explored through their thermodynamic signatures [14], using first-quantized parastatistics theories [8, 52, 53, 54] and para-manifolds [15], and as potential candidates for dark matter [16, 17, 18]. To date, the only experimental efforts to studying paraparticles are trapped ion simulations of para-bose and para-fermi oscillators [19, 20, 21, 22].
Recently, an alternative formulation was proposed which makes use of generalized commutation relations due to a four-tensor $R$ [23]. In this “$R$-paraparticle” formulation, quasiparticle excitations with parastatistical behavior can be constructed using nonlocal string operators from an underlying set of local spin operators. This approach evades the no-go theorems [10, 11] which assumes that excitations are only created by local operators. Moreover, the superselection rules [9] are inherently satisfied by quasiparticle excitations. This construction from local spin operators opens the possibility of realizing emergent paraparticles in condensed matter systems and in quantum simulations using trapped ions [24]. Furthermore, emergent $R$-paraparticles realized in higher dimensional condensed matter systems enable a novel way of nonlocal secret communication, as proposed in [6], which demonstrates their fundamental physical distinction from ordinary fermions and bosons, and provides an experimental way to observe their exotic exchange statistics. However, the discussions in [23] primarily focus on free paraparticle systems. In contrast, conventional bose and fermi systems exhibit a wealth of collective excitations and emergent phases arising from particle interactions. This naturally raises the question of how interactions might influence the behavior of paraparticles, making it an intriguing direction for further investigation.
@@
In this paper, we study a generalized version of the Luttinger model where we extend the conventional fermionic operators to $R$-paraparticle operators. This approach of extending the operators to $R$-paraparticle operators has recently been explored in the context of para-SYK models [36, 37]. We show that bosonization also occurs for $R$-paraparticles with fermi-like average occupation numbers [14], which we call “$R$-parafermions” in analogy to Green’s paraparticles [5, 9]. This presents a method in which interacting $R$-paraparticle systems can be probed using standard methods of solving bosonic systems [29]. Before proceeding to the remainder of the study, we first clarify that the term “parafermion” should not be confused with generalizations of Majorana fermions [38, 39, 40], which unfortunately share the same terminology in condensed matter literature. For instance, the work in [41] also addresses the bosonization of parafermions; however, the parafermions considered there are those related to Majorana fermions and are not the $R$-parafermions considered in this work.
The paper shall be organized as follows. In Sec. 2, we review the $R$-paraparticle formulation and clarify how $R$-parabosons and $R$-parafermions are classified based on their thermodynamics and exclusion statistics. We show that $R$-parafermions obey Pauli’s exclusion principle or one of its generalizations, prompting us to define $p$-orderedness in the $R$-paraparticle formalism. In Sec. 3, $R$-parafermionic density wave excitations are shown to be bosonic, independent of the number of degrees of freedom of the $R$-paraparticle, and are completely decoupled to flavor waves. However, for systems with two internal degrees of freedom, flavor waves, which are generalizations to spin waves, are bosonic only for specific $R$-parafermion species. More importantly, we show that bosonization only applies at low energies for $R$-paraparticles that satisfy Pauli exclusion principle, as $R$-parafermions obeying higher-order exclusion principles generate partition functions that do not match in the $R$-parafermionic and bosonic bases. These results point to possible experimental realizations of $R$-paraparticle signatures in one-dimensional systems by searching for signs of flavor-charge separation [31] in the absence of magnon excitations [42], as will be discussed in Sec. 4. Finally, we summarize the contents of this paper and provide possible extensions of this work in Sec. 5.
@@ 2.1 Formulation
where $x\equiv e^{-\beta\epsilon}$, $\beta=1/k_{B}T$, and $\epsilon$ is the $R$-paraparticle energy. $d_{n}$ are nonnegative integers and describe the dimension of the $n$-particle state. Importantly, $d_{n}=0$ implies that the $n$ $R$-paraparticles cannot occupy the same mode at once. The average occupation number then can be calculated as
@@ 2.2 R-parafermions
Analogous to Green’s formulation which generalizes parafermions with integer- and half-integer spins as parabosons and parafermions [5], we also construct a classification scheme for $R$-paraparticles. Suppose the GCR due to $R$ allows at most $n^{\prime}$ particles to occupy a single mode, irrespective of particle flavor. This implies that $d_{n>n^{\prime}}=0$, hence the average occupation number is reduced to
@@
i.e., the occupation number close to absolute zero is proportional to a Heaviside step function $\theta(x)$, and therefore the $R$-paraparticle possesses a fermi surface structure. This motivates us to classify $R$-paraparticles that satisfy Eq. (11b) as $R$-parafermions. Analogously, we show in the Appendix that when $d_{n}\neq 0$ for all $n$, the average occupation of the particle resembles bose statistics, and therefore these $R$-paraparticles are categorized as $R$-parabosons. In Green’s paraparticle formulation, these behaviors were also observed in [14].
The requirement that $d_{n>n^{\prime}}=0$ implies an exclusion principle is at play within the $R$-paraparticle’s GCR. In general, this can be expressed as
$(\hat{\psi}_{i,a}^{+})^{p}\neq 0\quad\text{while}\quad(\hat{\psi}_{i,a}^{+})^{p+1}=0,$ (12)
where $p=1$ is the usual Pauli exclusion principle. This definition is analogous to Green’s $p$-ordered parafermions, to which there can be at most $p$ particles in the same quantum state [9]. We therefore define a $p$-ordered $R$-parafermions as $R$-paraparticles that satisfy Eq. (12), i.e., allowing at most $p$ particles of the same flavor in a single mode. The maximum number of $R$-paraparticles that can occupy a single mode, $n^{\prime}$, need not be equal to $p$.
@@
where $\alpha\in\mathbb{C}_{\neq 0}$. The $R$-tensor of $M_{1}$ is exactly example 3 of [23] where $n^{\prime}=1$, while both $M_{2}$ and $M_{3}$ have $n^{\prime}=2$. Note that the matrices $M_{2}$ with $|\alpha|\neq 1$ and $M_{3}$ describe non-unitary $R$-tensors, and ordinary fermions are retrieved from $\alpha=-1$ in Eq. (13b). Likewise, $p=2$ cases also exist for $R$-paraparticles with $m=2$, such as those described by
@@
also with $\beta\in\mathbb{C}_{\neq 0}$. This also describes a non-unitary $R$-tensor whenever $|\beta|\neq 1$. The exclusion statistics of $R$-paraparticles described by $M_{4}$ go as $\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,b}^{+}=-\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,b}^{+}=0$ and $\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,a}^{+}\propto\hat{\psi}_{i,b}^{+}\hat{\psi}_{i,b}^{+}\neq 0$ for $\in\{1,2\}, aba\neq b$. This results to $\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,a}^{+}\propto\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,b}^{+}\hat{\psi}_{i,b}^{+}\allowbreak=0$, implying that three particles cannot occupy the same mode at once, and therefore $d_{n>2}=0$.
3 R-parafermion bosonization
Having defined $p$-ordering and the classification of $R$-paraparticles into $R$-parafermions or $R$-parabosons, let us now consider the bosonization of one-dimensional interacting $R$-paraparticles. We define the $R$-paraparticle Luttinger model kinetic Hamiltonian term as
@@ 3 R-parafermion bosonization
where the constant $\theta(rk-k_{F})$ imposes the normal ordering denoted by $\ldots:$, and $\pm k_{F}$ are the parafermi points of the 1D system. The indices $k$ and $a$ denote momentum and flavor, while $r\in\{+,-\}$ (a mode index) differentiates between the two branches of the dispersion relation $\varepsilon_{r}(k)=v_{F}(rk-k_{F})$, with $v_{F}$ as the fermi velocity. In the Luttinger model, the total density operator can be decomposed in terms of the left-moving and right-moving $R$-parafermions, $\hat{\rho}(k)=\hat{\rho}_{+}(k)+\hat{\rho}_{-}(k)$, where
@@ 3.1 Density wave bosonization and flavor-charge separation
which then imply that the density waves of all $R$-parafermionic Luttinger models have bosonic properties.
@@ 3.2 Flavor wave bosonization
$[\hat{\sigma}_{r}(-q),\hat{\sigma}_{r^{\prime}}(q^{\prime})]$ $=\sum_{k,k^{\prime},a,a^{\prime}}\alpha_{a}\alpha_{a^{\prime}}[\hat{\psi}^{+}_{r,k-q,a}\hat{\psi}_{r,k,a},\hat{\psi}^{+}_{r^{\prime},k^{\prime}+q^{\prime},a^{\prime}}\hat{\psi}_{r^{\prime},k^{\prime},a^{\prime}}]$
$=\sum_{k,k^{\prime},a,a^{\prime}}\alpha_{a}\alpha_{a^{\prime}}\Bigl{(}\hat{\psi}^{+}_{r,k-q,a}\hat{\psi}_{r,k,a}\hat{\psi}^{+}_{r^{\prime},k^{\prime}+q^{\prime},a^{\prime}}\hat{\psi}_{r^{\prime},k^{\prime},a^{\prime}}$
$\quad-\hat{\psi}^{+}_{r^{\prime},k^{\prime}+q^{\prime},a^{\prime}}\hat{\psi}_{r^{\prime},k^{\prime},a^{\prime}}\hat{\psi}^{+}_{r,k-q,a}\hat{\psi}_{r,k,a}\Bigr{)}$
$=\sum_{k,k^{\prime},a,a^{\prime}}\alpha_{a}\alpha_{a^{\prime}}\biggl{\{}\delta_{r,r^{\prime}}\delta_{k-q,k^{\prime}}\delta_{a,a^{\prime}}\hat{\psi}^{+}_{r,k-q,a}\hat{\psi}_{r^{\prime},k^{\prime},a^{\prime}}$
$\qquad+\biggl{(}\sum_{b,c,d,e,f,g}R^{ab}_{a^{\prime}c}R^{de}_{ab}R^{a^{\prime}c}_{gf}\hat{\psi}^{+}_{r^{\prime},k^{\prime}+q^{\prime},d}\hat{\psi}^{+}_{r,k-q,e}\hat{\psi}_{r^{\prime},k^{\prime},f}\hat{\psi}_{r,k,g}\biggr{)}$
$\qquad-\hat{\psi}^{+}_{r^{\prime},k^{\prime}+q^{\prime},a^{\prime}}\hat{\psi}_{r^{\prime},k^{\prime},a^{\prime}}\hat{\psi}^{+}_{r,k-q,a}\hat{\psi}_{r,k,a}\biggr{\}}$ (20)
explicitly for specific $R$-tensors. In this case, bosonic flavor waves occur only when the RHS of Eq. (3.2) reduces to the RHS of Eq. (16). This requirement is equivalent to
@@
as the second term in Eq. (21) cancels out the last term in Eq. (3.2) and the remaining Kronecker delta terms reduce to $\hat{e}_{ij}$. When the above requirement is satisfied, Eq. (3.2) becomes
@@
allows for the establishment of $\hat{c}_{k},\hat{c}^{+}_{k}]=\delta_{kk^{\prime}}$. $R$-tensors in the form of Eq. (13a), (13b), and (13d) with $\beta^{2}=1$ all follow these results. However, $R$-parafermions described by Eqs. (13c) and (13d) with $\beta^{2}\neq 1$ do not satisfy Eq. (21) and therefore their $R$-parafermionic Luttinger models do not have bosonic flavor wave profiles.
@@ 3.3 Equivalence spectrum
We now check wether the system described using $R$-parafermion operators are the same as that described by bosonic ones. For $m=2$ $R$-parafermions, it can be shown that the free Luttinger model Hamiltonian obeys the commutation relation
@@
where $\hat{n}_{k}=\hat{b}_{k}^{+}\hat{b}_{k}$ and $\hat{N}_{r,a}=\sum_{k}(\hat{\psi}_{r,k,a}^{+}\hat{\psi}_{r,k,a}-\theta(rk-k_{F}))$. Note that these results are applicable only to $R$-parafermions with bosonic flavor waves. The second summation term takes care of the normal ordering imposed in the $R$-parafermion operator basis. The flavor wave contribution to the Hamiltonian is the same as Eq. (25), but with $\hat{b}_{k}^{+}\hat{b}_{k}$ replaced by $\hat{c}_{k}^{+}\hat{c}_{k}$ [29], hence we consider the total Hamiltonian
@@
where $H_{0}$ is denoted by Eqs. (14) and (25) in the $R$-parafermion and boson operator bases, respectively.
One way to verify the correspondence between these Hamiltonian forms is to show that their degeneracies match in both bases. The case where $R$-parafermions have no internal degrees of freedom (i.e., spinless fermions) has already been done in [26] by evaluating the partition functions of each basis separately. Following this method, the $R$-parafermion partition function is calculated as
@@
where $y\equiv e^{-\beta v_{F}\pi/L}$. The (2n1)(2n-1) exponent is obtained by assuming that the system is under periodic boundary conditions such that $k=2n\pi/L$ (where $n$ are energy level indices) and $k_{F}=n_{F}\pi/L$. The variable $p$ denotes the order of the $R$-parafermion as discussed in Sec. 2.2. In particular, for $p=1$, we have
@@
Comparison between Eqs. (28a) and (29) shows that in the low-temperature limit (i.e., $\beta\to\infty$, and subsequently $y\to 0$), the boson and $p=1$ $R$-parafermion partition functions coincide, hence the spectrum of the free Luttinger model is equivalent in the $R$-paraparticle and boson operator basis only if the $R$-parafermion obeys Pauli’s exclusion principle. This result generalizes the idea that the Luttinger model is acceptable only at low-energies [46]. On the other hand, if the $R$-parafermion possesses a higher-order exclusion principle (such as in the case of $p=2$ in Eq. (28b)), $z_{PF}$ quickly diverges. This means that bosonization is not applicable for $R$-parafermions with $p>1$, and possibly the Luttinger model is not appropriate for modelling such systems.